Right triangle: side a =x, side b = 3x, hyp. = 19 what is x?
Thanks,
Dave

Answer by scott8148(6628)

Pythagoras says ___ x^2 + (3x)^2 = 19^2

10x^2 = 361

x = sqrt(361 / 10)